The BRS Corporation makes collections on sales according to the following schedule:

45% in month of sale
48% in month following sale
7% in second month following sale

The following sales have been budgeted:

Sales
April $ 120,000
May $ 100,000
June $ 140,000

Budgeted cash collections in June would be:

Multiple Choice

  • $119,400

  • $140,000

  • $140,840

  • $111,000
Answer #1

Computation of Budgeted cash collection in the month of June:

Collection of cash in the same month = $140000*45% = $63000

Collection of May month sale in June = $100000*48% = $48000

Collection of April month sale in June = $120000*7% = $8400

Total cash collected in the month of June = $63000 + $48000 + $8400 = $119400.